# Revealing Language Model Trajectories via Kullback-Leibler Divergence

### v1

@@ Abstract
A recently proposed method enables efficient estimation of the KL divergence between language models, including models with different architectures, by assigning coordinates based on log-likelihood vectors. To better understand the behavior of this metric, we systematically evaluate KL divergence across a wide range of conditions using publicly available language models. Our analysis covers comparisons between pretraining checkpoints, fine-tuned and base models, and layers via the logit lens. We find that trajectories of language models, as measured by KL divergence, exhibit a spiral structure during pretraining and thread-like progressions across layers. Furthermore, we show that, in terms of diffusion exponents, model trajectories in the log-likelihood space are more constrained than those in weight space.
Revealing Language Model Trajectories via Kullback-Leibler Divergence
Ryo Kishino1  Yusuke Takase1  Momose Oyama1,2 Hiroaki Yamagiwa1   Hidetoshi Shimodaira1,2 1Kyoto University 2RIKEN kishino.ryo.32s@st.kyoto-u.ac.jp, {y.takase, oyama.momose}@sys.i.kyoto-u.ac.jp, {h.yamagiwa, shimo}@i.kyoto-u.ac.jp
@@ 1 Introduction
Figure 1: Visualization of the training trajectories of Pythia models using t-SNE applied to log-likelihood vectors. Each point represents a model checkpoint, with lighter colors indicating earlier steps. The thickness of the connecting lines reflects the magnitude of KL divergence between checkpoints. The top panel shows trajectories for five model sizes and multiple random seeds. The bottom panel shows only the five sizes, with both 2D and 3D t-SNE visualizations for clarity. See Section 3 for details.
@@
(a) KL divergence for Pythia 410M, measured between checkpoints saved during pretraining, different random seeds, and different model sizes. For the inter-size comparison, the final checkpoints were used. See Section 3 for details.
Type (#pair) Pair Median Mean SD
mistral (25) fine-tuning 0.32 0.37 0.32
random 0.79 0.74 0.39
llama-3 (20) fine-tuning 0.20 0.67 0.80
random 0.66 0.87 0.66
all (66) fine-tuning 0.40 1.5 3.6
random 2.1 5.0 9.5
(b) Summary statistics (median, mean, SD) of KL divergence between fine-tuned and base models, and between random model pairs. See Section 4 for details.
Model 1$\to$2 16$\to$17 31$\to$32
@@
(c) KL divergence between adjacent layers within models, measured using the logit lens. See Section 5 for details.
@@
The growing diversity of language models has created a need to quantify their differences systematically. Treating them as probability distributions over text, the Kullback-Leibler (KL) divergence serves as a natural comparison metric. Oyama et al. (2025) proposed using log-likelihood vectors on a fixed text set to assign model coordinates in a Euclidean space. In this space, the squared distance between models approximates their KL divergence, making model comparison a geometric problem.
However, their analysis did not cover the typical KL divergence values. As an early report on model analysis leveraging log-likelihood vectors, we use publicly available models to measure the KL divergence between checkpoints saved during pretraining (Section 3), fine-tuned and base models (Section 4), and different layers (Section 5) as shown in Table 1.
Furthermore, we analyze the trajectories of language models throughout training and across layers, yielding several notable observations. As shown in Fig. 1, portions of the pretraining trajectories exhibit a spiral structure. Moreover, from the perspective of anomalous diffusion, model dynamics in the log-likelihood space become increasingly constrained during the later steps of training, compared to those in weight space.
2 Review: Model Map
Log-likelihood vector.
The probability that a language model $p$ generates a text (i.e., a sequence of tokens) $x=(y_{0},\ldots,y_{M})$ is given by $p(x)=\prod_{m=1}^{M}p(y_{m}\mid y_{0},\ldots,y_{m-1})$. Oyama et al. (2025) defined the log-likelihood vector of a language model $p$ over a predefined text set $\{x_{1},\ldots,x_{N}\}$ as
$\bm{\ell}=(\log p(x_{1}),\ldots,\log p(x_{N}))^{\top}\in\mathbb{R}^{N},$
and showed that it serves as a useful feature representation for comparing models and predicting downstream task performance.
KL divergence.
Let $\{p_{i}\}_{i=1}^{K}$ be a set of language models, and $\bm{L}=(\bm{\ell}_{1},\ldots,\bm{\ell}_{K})^{\top}\in\mathbb{R}^{K\times N}$ be the matrix formed by stacking the log-likelihood vectors $\bm{\ell}_{i}$ of each model. Define $\bm{Q}=(\bm{q}_{1},\ldots,\bm{q}_{K})^{\top}$ as the matrix obtained by double-centering [^1] Row-wise centering followed by column-wise centering. $\bm{L}$. Then, the KL divergence between two models $p_{i}$ and $p_{j}$ can be approximated as
$2\mathrm{KL}(p_{i},p_{j})\approx\|\bm{q}_{i}-\bm{q}_{j}\|^{2}/N.$
In other words, by associating each model $p_{i}$ with a coordinate $\bm{q}_{i}\in\mathbb{R}^{N}$, KL divergence is approximated by squared Euclidean distance in this space. This coordinate system is referred to as the model map in Oyama et al. (2025). In particular, we refer to a sequence of models represented as coordinates in the model map as a trajectory.
General experimental settings.
For the text set, we use the same 10,000 texts extracted from the Pile Gao et al. (2020) as in Oyama et al. (2025) [^2] https://github.com/shimo-lab/modelmap/. The values of KL divergence are normalized by the average text length, $\bar{B}=972.3188$ bytes, and are reported in units of bits per byte. In particular, the matrix $\bm{Q}$ is divided by $\sqrt{2N\bar{B}\log 2}$ in advance, so that the squared Euclidean distance directly approximates the KL divergence in bits per byte. All visualizations are based on this $\bm{Q}$.
3 Pretraining
In the Pythia suite Biderman et al. (2023), the checkpoints saved during pretraining are publicly available for each model size. Additionally, van der Wal et al. (2025) released checkpoints for smaller Pythia models trained with different random seeds. In this section, we investigate how a language model’s trajectory evolves during pretraining by analyzing these checkpoints.
Settings.
We use checkpoints from the standard Pythia models with sizes 410M, 1B, 1.4B, 2.8B, and 6.9B. In addition, for the 410M model, we supplement our analysis with checkpoints from seven different random seeds. The checkpoints are saved at training steps 0, 1, 2, 4, …, 512, 1k, 2k, 3k, …, and 143k. We excluded the 1B checkpoint at step 116k due to anomalous weights, and removed texts with irregular log-likelihood variations across checkpoints. See Appendix D.1 for details.
KL divergence.
Table 1(a) shows the KL divergence for Pythia 410M. The KL divergence between checkpoints saved every 1k steps gradually decreases as training progresses, and after step 10k, it stays within a small range (approximately 0.02 to 0.07 bit/byte), as also shown in the right panel of Fig. 4. Moreover, the values between different random seeds are around 0.1 bit/byte, while those between different sizes range from 0.3 to 2 bit/byte, indicating larger differences. See Appendix D.2 for details.
Trajectory.
Figure 2: Trajectory of Pythia 2.8B during pretraining, visualized by applying 3D t-SNE, followed by PCA on the later half portion. (Left) PC2 vs. PC3. (Right) Autocorrelation function of PC3.
Model Size 410M 1B 1.4B 2.8B 6.9B
Period (step) 30k 13k 15k 14k 12k
Table 2: Period of the spiral trajectory of Pythia during pretraining. See Appendix D.3 for details.
Fig. 1 shows a simultaneous visualization of the pretraining trajectories of five model sizes using t-SNE van der Maaten and Hinton (2008) [^3] We used the scikit-learn implementation with the default perplexity value of 30., which is not possible with weight-based representations. The trajectories evolve continuously during training [^4] The apparent jumps in the 2D trajectories of the 2.8B and 6.9B models are artifacts of t-SNE, which tends to separate overlapping points. These trajectories appear continuous in 3D t-SNE visualizations.. Moreover, the 3D visualization suggests that the trajectories exhibit a spiral structure. As shown in Fig. 2, applying PCA to the later portion of the trajectory obtained via t-SNE reveals a circular pattern. The autocorrelation function of PC3 exhibits periodicity. The estimated periods of the spiral trajectories are shown in Table 2. Similar phenomena in weight trajectories have been studied previously Cohen et al. (2021); Netay (2024) described in detail in Appendix A. In addition, refer to Appendix D.4 for a trajectory visualization using PCA.
Comparison with weights.
Figure 3: Temporal evolution of the squared Euclidean distance in the weights and log-likelihood vectors from step 120k to step 130k for Pythia. The diffusion exponent $c$ is estimated using the least squares method.
Figure 4: (Left) Diffusion exponent $c$ as a function of the starting point $t_{0}$. (Right) KL divergence between consecutive checkpoints of Pythia (corresponding to Table 1(a)). See Appendix D.5 for details.
Differences between models with the same architecture can be measured by the distance between their weights. Let $\bm{W}_{t}$ be the weights of a neural network at step $t$. According to Kunin et al. (2024), after model performance has converged, the weights diffuse following a power law with exponent $c$ [^5] Brownian motion corresponds to $c=1$, while anomalous diffusion is defined as $c\neq 1$. Specifically, $c>1$ means superdiffusion, and $c<1$ means subdiffusion. If the evolution is fractional Brownian motion Mandelbrot and Van Ness (1968), the Hurst exponent corresponds to $H=c/2$.. That is, with respect to an initial step $t_{0}$, the relationship $\|\bm{W}_{t}-\bm{W}_{t_{0}}\|^{2}\propto|t-t_{0}|^{c}$ holds, where $c$ may differ from $1$.
To verify this, we measured the evolution of the squared Euclidean distance in both the weights $\bm{W}$ and the log-likelihood vectors $\bm{q}$, over steps 120k to 130k of Pythia. As shown in Fig. 3, both $\bm{W}$ and $\bm{q}$ follow a power-law relationship with exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ respectively. The weights $\bm{W}$ exhibit Brownian-like diffusion ($c_{\mathrm{w}}\approx 1$), whereas the diffusion of $\bm{q}$ is substantially suppressed ($c_{\mathrm{q}}\approx 0.2$). In addition, Fig. 4 illustrates how the exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ vary with the starting step. Notably, $c_{\mathrm{q}}$ remains consistently smaller than $c_{\mathrm{w}}$ throughout. Both $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ tend to decrease as training progresses. However, around step 120k, when the KL divergence between successive checkpoints stabilizes, $c_{\mathrm{w}}$ increases, while $c_{\mathrm{q}}$ shows no significant change. As discussed in Appendix B, the difference in dynamics between these two spaces could be explained by the flat minima hypothesis Keskar et al. (2017) and by constraints imposed by the loss function on the log-likelihood space.
4 Fine-tuning
Figure 5: Map showing base models and their fine-tuned variants. Lines represent parent-child relationships between models. Colors indicate model types as defined in Oyama et al. (2025). Generation refers to the depth from the root model within each lineage.
The introduction of efficient fine-tuning methods Hu et al. (2022); Rafailov et al. (2023) has led to the development of a large number of fine-tuned models. In this section, we investigate the KL divergence between fine-tuned and base models.
Settings.
Out of the 1,018 language models analyzed in Oyama et al. (2025), we use 66 pairs of fine-tuned and base models (87 models in total; Fig. 5). The parent-child relationships are identified using Hugging Face Hub API. See Appendix E for details.
KL divergence.
Table 1(b) shows the statistics on the KL divergence between fine-tuned and base models, and between randomly selected models. For the cumulative distribution, see Fig. 18 in Appendix E. KL divergence values tend to be slightly smaller for models of type mistral or llama-3. In all cases, the KL divergence between fine-tuned and base models is smaller than that between random pairs, with a median ranging from 0.2 to 0.4 bit/byte.
5 Layer
Figure 6: t-SNE visualization of trajectories across layers for 50 models, obtained by applying the logit lens to each layer. Deeper layers are represented with lighter colors, and the final layer is indicated by a white square.
Figure 7: For 10 models per model type: (left) t-SNE visualization of trajectories across layers, where the final layer is indicated by a white square; (right) KL divergence between adjacent layers (corresponding to Table 1(c)). See Fig. 20 in Appendix F for a 3D plot.
We investigate model trajectories across layers by treating the network up to each layer as an individual model using the logit lens nostalgebraist (2020).
Settings.
We use a subset of the 1,018 language models analyzed in Oyama et al. (2025). First, without restricting model types, we select the 50 most downloaded models (Fig. 6). When limiting to specific model types, we use the 10 most downloaded models from each of the llama-2, llama-3, and mistral with 32 layers (30 models in total; Fig. 7). See Appendix F for details on the models used.
KL divergence.
The right panel of Fig. 7 and Table 1(c) show the KL divergence between consecutive layers. Models of the same type exhibit similar trends. For llama-2, Wendler et al. (2024) reported a sharp drop in the entropy of next-token distributions at intermediate layers in a translation task, which corresponds to the peak positions in Fig. 7. The KL divergence between consecutive layers ranges from as low as 1 bit/byte to as high as 10 to 25 bit/byte, indicating substantial variation.
Trajectory.
Fig. 6 and the left panel of Fig. 7 illustrate the model trajectories across layers. The trajectories exhibit jumps at KL divergence peaks and tend to align in deeper layers. The trajectories align with increasing mean log-likelihood (Fig. 19, Appendix F).
6 Conclusion
We analyzed KL divergence and trajectories across checkpoints, model pairs, and layers, revealing spiral and aligned structures over training and depth.
Limitations
  • The checkpoints for Pythia are available only at 1k-step intervals except for the early stages of training. When checkpoints are saved at shorter intervals, the resulting trajectories can differ from our observation. In particular, the estimation of the diffusion exponent may be affected by errors introduced by the coarse granularity of the 1k-step resolution.
  • The KL divergence we measured depends on the text set used. In future work, we plan to investigate how the KL divergence and model trajectories vary when a different text set is used.
  • The number of models used is limited in the experiment measuring KL divergence between fine-tuned and base models. This is because some fine-tuned models do not include information about their base models in their model cards obtained by using Hugging Face Hub API.
  • In the layer-wise analysis, the logits obtained from shallow layers via the logit lens contain substantial noise. This issue could be addressed by using the tuned lens Belrose et al. (2023), an improved version of the logit lens. However, the number of publicly available pretrained tuned lenses is currently limited. We believe this limitation can be overcome in the future by training tuned lenses ourselves.
  • While we observe structured behaviors such as spiral trajectories and subdiffusion in the log-likelihood space, their underlying causes and implications for learning dynamics have not been quantitatively analyzed and remain open for future investigation.
Acknowledgments
We thank Luis Iván Hernández Ruíz for helpful discussions. This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST BOOST JPMJBS2407.
Appendix A Related Work on Oscillatory Trajectories in Weight Space
We analyzed model trajectories in the log-likelihood space, where each model was represented by its output probabilities on a fixed text set. This space reflects the model’s behavior directly, and our study was, to the best of our knowledge, the first to reveal structural patterns in training dynamics within this representation. In particular, we observed the emergence of spiral-shaped trajectories as well as indications of constrained diffusion, which we further examine in Appendix B.
While our main focus was on the log-likelihood space, we also briefly examined model trajectories in weight space for comparison (Appendix D.5). As shown in Fig. 16, oscillatory behavior was also observed in the weight space during the later stages of training. Although the patterns differed in periodicity and structure, the presence of oscillations in both spaces suggests a potentially broader phenomenon in the dynamics of neural networks. The observed difference may reflect the presence of multiple modes of variation, with different modes being emphasized in the weight and output spaces.
To contextualize these findings, we review prior work that has reported oscillatory behaviors in weight space during training. The Edge of Stability phenomenon Cohen et al. (2021); Arora et al. (2022) refers to the behavior of training loss oscillating while decreasing, observed when gradient descent enters an unstable region where the largest Hessian eigenvalue hovers near the critical learning rate. Singh et al. (2025) observed micro-level oscillations in training trajectories through changes in the cosine similarity of weights. Netay (2024) empirically identified double-helical trajectories in neural network weights, noting a relationship between these trajectories’ periodicity and the hyperparameters of the Adam optimizer. Kunin et al. (2024) found that weights exhibit anomalous diffusion after convergence and demonstrated oscillatory dynamics in the weight phase space, governed by the eigenstructure of the Hessian of the loss function.

### v2

# Establishing a Scale for Kullback–Leibler Divergence in Language Models Across Various Settings

@@ Abstract
Log-likelihood vectors define a common space for comparing language models as probability distributions, enabling unified comparisons across heterogeneous settings. We extend this framework to training checkpoints and intermediate layers, and establish a consistent scale for KL divergence across pretraining, model size, random seeds, quantization, fine-tuning, and layers. Analysis of Pythia pretraining trajectories further shows that changes in log-likelihood space are much smaller than in weight space, resulting in subdiffusive learning trajectories and early stabilization of language-model behavior despite weight drift.
Establishing a Scale for Kullback–Leibler Divergence
in Language Models Across Various Settings
Ryo Kishino1   Yusuke Takase1   Momose Oyama1,2 Hiroaki Yamagiwa1   Hidetoshi Shimodaira1,2 1Kyoto University 2RIKEN kishino.ryo.32s@st.kyoto-u.ac.jp, {oyama.momose, y.takase}@sys.i.kyoto-u.ac.jp, {h.yamagiwa, shimo}@i.kyoto-u.ac.jp
@@ 1 Introduction
To understand learning dynamics and intermediate-layer representations in language models, it is essential to quantify changes in model behavior and compare them across models. While traditional analyses rely on weight parameters Chen et al. (2022); Kunin et al. (2024); Singh et al. (2025); Gigante et al. (2019), weight permutation symmetries Hecht-Nielsen (1990) and architectural dependencies hinder direct comparisons between models with different learning methods or designs.
@@
Figure 1: Visualization of the pretraining trajectories of Pythia models across five model sizes and seven random seeds using t-SNE (perplexity=30) applied to log-likelihood vectors. Each point represents a model checkpoint, with lighter colors indicating earlier steps. The thickness of the connecting lines reflects the magnitude of KL divergence between successive checkpoints. See Appendix D for additional visual analyses.
Oyama et al. (2025) proposed the log-likelihood vector, a common representation of a model’s probabilities over a set of texts. They showed that models with different architectures can be embedded in the same space. Moreover, since the squared Euclidean distance in this space estimates the Kullback–Leibler (KL) divergence, model comparison can be formulated as a geometric problem.
In this study, we demonstrate that log-likelihood vectors provide a consistent representation not only for fully trained models but also for training checkpoints, quantized models, and intermediate layers. We therefore handle diverse models within a single coordinate system (see Figs. 1 and 2). We establish KL divergence scales across various settings, including checkpoints, model sizes, random seeds, quantization, fine-tuning, and layers as practical metrics for model comparison.
Figure 2: (a) t-SNE visualization (perplexity=30) of the original models and their 8-bit quantized models. (b) t-SNE visualization (perplexity=20) of the base models and their fine-tuned models. Lines represent parent-child relationships. Generation refers to the depth from the root model within each lineage. (c) t-SNE visualization (perplexity=30) of the trajectories across layers for 50 models, obtained by applying the logit lens to each layer. Deeper layers are represented with lighter colors, and the final layer is indicated by a white square.
Moreover, experiments on Pythia show that when focusing on pretraining trajectories and estimating KL-scaling exponents, large changes in weight space remain remarkably small in log-likelihood space. This discrepancy appears as subdiffusive behavior in later training steps, indicating that the model’s output is far more stable.
2 Preliminaries: Model Map and KL
Log-likelihood vector.
The probability that a language model $p$ generates a text (i.e., a sequence of tokens) $x=(y_{0},\ldots,y_{M})$ is given by $p(x)=\prod_{m=1}^{M}p(y_{m}\mid y_{0},\ldots,y_{m-1})$. Oyama et al. (2025) defined the log-likelihood vector of a model $p$ over a predefined text set [^1] Olsson et al. (2022) used token-level log-likelihoods. $\{x_{1},\ldots,x_{N}\}$ as
$\bm{\ell}=(\log p(x_{1}),\ldots,\log p(x_{N}))^{\top}\in\mathbb{R}^{N},$
and showed that it is a useful feature representation for model comparison and downstream performance prediction. This method has also been used in subsequent work such as Harada et al. (2025).
KL divergence.
Let $\{p_{i}\}_{i=1}^{K}$ be a set of language models, and $\bm{L}=(\bm{\ell}_{1},\ldots,\bm{\ell}_{K})^{\top}\in\mathbb{R}^{K\times N}$ be the matrix formed by stacking the log-likelihood vectors $\bm{\ell}_{i}$ of each model. Define $\bm{Q}=(\bm{q}_{1},\ldots,\bm{q}_{K})^{\top}$ as the matrix obtained by double-centering [^2] Row-wise centering (over texts) and column-wise centering (over models), respectively, of the log-likelihood matrix. $\bm{L}$. Then, the KL divergence between two models $p_{i}$ and $p_{j}$ can be approximated as
$2\mathrm{KL}(p_{i},p_{j})\approx\|\bm{q}_{i}-\bm{q}_{j}\|^{2}/N.$
In other words, by associating each model $p_{i}$ with a coordinate $\bm{q}_{i}\in\mathbb{R}^{N}$, KL divergence is approximated by squared Euclidean distance in this space. Oyama et al. (2025) referred to this coordinate system as the model map. For the interpretation of the KL divergence, the standard error, and the dependence on the text set, see Appendix C.
Model map for various settings.
Log-likelihood vectors can be computed not only for fully trained language models but also for models during training (Figs. 1 and 2b) and for quantized models (Fig. 2a). Moreover, by treating the network up to each layer as a single model via the logit lens nostalgebraist (2020), intermediate layers can also be analyzed (Fig. 2c). These analyses are difficult to carry out with weight-based approaches Gigante et al. (2019); Chen et al. (2022); Kunin et al. (2024); Singh et al. (2025).
General experimental settings.
For the text set, we use the same 10,000 texts extracted from the Pile arxiv:2101.00027 as in Oyama et al. (2025). The values of KL divergence are normalized by the average text length, $\bar{B}=972.3188$ bytes, and are reported in units of bits per byte. Accordingly, the matrix $\bm{Q}$ is divided in advance by $\sqrt{2N\bar{B}\log 2}$, so that the squared Euclidean distance between rows of $\bm{Q}$ directly approximates the KL divergence in bits per byte.
Figure 3: Scale of KL divergence (bits per byte) across various settings. The first four entries from the left correspond to pretraining-related comparisons, showing the KL divergence between consecutive saved checkpoints at 1k-step intervals during the last 50k steps of training, at 1k-step intervals during the first 50k steps, across different random seeds at the final checkpoint, and across different model sizes. The subsequent entries show the KL divergence before and after quantization (8-bit and 4-bit), before and after fine-tuning, between randomly selected pairs within the same model type, between randomly selected pairs across all model types, and between adjacent layers. The vertical axis is shown on a symlog scale. See Appendix A and Table 1 for more details.
3 KL Divergence Across Various Settings
In this section, we measure and summarize KL divergence under a variety of settings, and present the results in Fig. 3. KL divergence is an interpretable quantity on an absolute scale; for example, a value of 0.1 bits/byte corresponds to only 0.1 bit per 8-bit text unit, whereas 10 bits/byte indicates a very large difference, exceeding the 8-bit scale of the text itself [^3] Since KL divergence is unbounded, it can exceed 8 bits/byte despite the 8-bit representation of text.. As a reference point, the effective per-byte entropy of the text set in our experiments is approximately 0.71 bits/byte (see Appendix C). Our experiments reveal that each setting is associated with a characteristic scale of KL divergence. Even the same value of 0.1 bits/byte can be interpreted as a substantial change between consecutive saved 1k-step checkpoints during training, but as a minor difference when comparing models of different types.
3.1 Pretraining
Settings.
We use publicly available pretraining checkpoints from the Pythia suite arxiv:2304.01373 for model sizes 410M, 1B, 1.4B, 2.8B, and 6.9B. For the 410M model, checkpoints are available for nine random seeds (1–9) released by wal2025polypythias. From these, we exclude seeds 3 and 4 due to training loss spikes detected via KL divergence, leaving seven seeds for our experiments. Checkpoints are saved at training steps 0, 1, 2, 4, …, 512, 1k, 2k, 3k, …, and 143k. We exclude the 1B checkpoint at step 116k due to anomalous weights and remove texts exhibiting irregular log-likelihood variations across checkpoints. Details of the preprocessing, data filtering, and anomaly detection procedures for pretraining are provided in Appendix E.
KL divergence.
For all model sizes, the KL divergence between consecutive saved checkpoints at 1k-step intervals after the warmup phase is typically 0.05–0.1 bits/byte during the early stage of training, and becomes extremely small in the later stage, ranging from about 0.01 to 0.05 bits/byte. At the final checkpoint, the KL divergence across different random seeds for the 410M model is around 0.1 bits/byte, while that across different model sizes ranges from 0.15 to 1.7 bits/byte, indicating substantially larger differences.
3.2 Quantization
Settings.
We analyze the impact of post-training quantization Dettmers et al. (2022). We use a subset of the 1,018 language models analyzed in Oyama et al. (2025), selecting the 50 most downloaded models (Fig. 2a); see Appendix F for details on the models used.
KL divergence.
The median KL divergence before and after 8-bit quantization is 0.44 bits/byte, while it is slightly larger for 4-bit quantization at 0.49 bits/byte. As shown in Fig. 3, the distributions exhibit low variance in both cases, indicating that the degree of degradation induced by quantization is relatively consistent across models. Furthermore, in Fig. 2a, the shifts induced by quantization appear to be aligned in both direction and magnitude within each model type, suggesting that quantization acts as a structured perturbation in the log-likelihood space rather than as random noise.
3.3 Fine-tuning Effects
Settings.
Out of the 1,018 language models analyzed in Oyama et al. (2025), we use 66 pairs of fine-tuned and base models (87 models in total; Fig. 2b). The parent-child relationships are identified using Hugging Face Hub API. See Appendix G for details.
KL divergence.
Fig. 3 shows that fine-tuning induces a relatively small but non-negligible change within a model type, smaller than the variation among randomly paired models of the same type and much smaller than the variation across model types; the corresponding median KL divergence values are 0.40, 0.95, and 2.2 bits/byte.
3.4 Across Layers
Settings.
We investigate model trajectories across layers by treating the network up to each layer as an individual model using the logit lens nostalgebraist (2020). We use the same set of 50 models as those used in Section 3.2 (Fig. 2c).
KL divergence.
The KL divergence between adjacent layers spans a wide range, with a median of 3.0 bits/byte and a standard deviation of 13 bits/byte. Moreover, as shown in Fig. 2c, models within the same type exhibit similar trajectory trends across layers.
4 Scaling Analysis of KL Divergence Along Pretraining Trajectories
In this section, we examine the scaling behavior of changes in KL divergence over training time along pretraining trajectories of Pythia models, using the same checkpoints as in Section 3.1.
4.1 Trajectory on the Log-Likelihood Space
Figure 4: KL divergence between consecutive saved checkpoints of Pythia during pretraining. (Left) Warmup phase with non-uniform checkpoint spacing. (Right) Post-warmup phase with checkpoints saved every 1k training steps.
Fig. 1 visualizes pretraining trajectories of Pythia models across five model sizes and seven random seeds in the log-likelihood space using t-SNE, enabling joint comparisons that are difficult in weight space [^4] Two-dimensional t-SNE visualizations may exhibit apparent jumps due to dimensionality constraints, whereas three-dimensional t-SNE visualizations confirm continuity in the original high-dimensional space; see Appendix D.. Fig. 4 shows the KL divergence between consecutive saved checkpoints during pretraining. After the warmup phase, the KL divergence between checkpoints saved every 1k steps decreases and, after step 10k, remains within a narrow range regardless of model size, whereas substantially larger values are observed during warmup.
4.2 Comparison with Weight Space
We compare the stability of training trajectories by estimating diffusion exponents that characterize the power-law scaling of KL divergence. According to Kunin et al. (2024), after model performance has converged, the weights diffuse following a power law with exponent $c_{\mathrm{w}}$. Specifically, with respect to an initial step $t_{0}$, the squared Euclidean distance satisfies $\|\bm{W}_{t}-\bm{W}_{t_{0}}\|^{2}\propto|t-t_{0}|^{c_{\mathrm{w}}}$. We show that an analogous power-law relationship also holds for the log-likelihood vectors $\bm{q}_{t}$, namely $\|\bm{q}_{t}-\bm{q}_{t_{0}}\|^{2}\propto|t-t_{0}|^{c_{\mathrm{q}}}$, which corresponds to the KL divergence.
In Fig. 5, the weights $\bm{W}$ exhibit Brownian-like diffusion with $c_{\mathrm{w}}\approx 1$, whereas the diffusion of $\bm{q}$ is substantially suppressed with $c_{\mathrm{q}}\approx 0.2$. Under a fractional Brownian motion interpretation, these exponents correspond to effective fractal dimensions of $D_{\mathrm{w}}\approx 2$ and $D_{\mathrm{q}}\approx 10$, respectively [^5] For diffusion exponent $c$, Brownian motion corresponds to $c=1$, while $c\neq 1$ indicates anomalous diffusion. If the process is fractional Brownian motion Mandelbrot and Van Ness (1968), the Hurst exponent is $H=c/2$, and the Hausdorff dimension of the trajectory is $D=1/H$ (Adler, 1981, Theorem 8.4.1).. In addition, the right panel of Fig. 5 illustrates how the exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ vary with the starting step $t_{0}$. Compared to weight trajectories, the scaling of KL divergence in log-likelihood space exhibits greater geometric complexity due to many-to-one folding, thus remaining confined to a narrow region and suggesting early stabilization of language-model behavior despite continued drift in weight space.
The difference in dynamics between these two spaces may be attributed to folding induced by permutation symmetries in multilayer neural networks Fukumizu and Amari (2000) and to the flat minima hypothesis Keskar et al. (2017). A detailed discussion is provided in Appendix B.
Figure 5: (Left) Temporal evolution of the squared Euclidean distance in the weights and log-likelihood vectors from steps 120k to 130k for Pythia. (Right) Diffusion exponent as a function of the starting point $t_{0}$. The exponent is estimated using the least squares method.
5 Conclusion
We measured KL divergence between language models across a wide range of conditions and established a unified and interpretable scale. We further examined the scaling behavior of changes in KL divergence along learning trajectories.
Limitations
  • We conduct our experiments using the same set of 10,000 text chunks as those used in Oyama et al. (2025), and the measured KL divergence therefore depends on the choice of the text set. However, we verified that different random subsets of 10,000 text chunks from the same Pile corpus yield highly consistent KL estimates (Appendix C), while the impact of domain shifts across text sets has not been examined.
  • The set of models analyzed in this study is limited and does not fully cover the diversity of existing language models. In particular, pretraining trajectory analyses are conducted only on the Pythia family, and experiments involving fine-tuning, quantization, and layer-wise analysis rely on selected subsets of models due to data availability and metadata constraints, such as incomplete base-model information in model cards.
  • In the layer-wise analysis, the logits obtained from shallow layers via the logit lens contain substantial noise. This issue could be addressed by using the tuned lens Belrose et al. (2023), an improved version of the logit lens. However, the number of publicly available pretrained tuned lenses is currently limited. We believe this limitation can be overcome in future work by training tuned lenses.
  • The checkpoints for Pythia are available only at 1k-step intervals except for the early stages of training. Accordingly, our analysis focuses on trajectory dynamics at scales of 1k steps or larger, while finer-scale behavior below this resolution is not examined. This limitation does not affect our analysis of dynamics at coarser temporal scales.
  • While we observe structured behaviors such as subdiffusion in the log-likelihood space, this work represents an initial attempt at trajectory analysis in this space, and our analysis of their underlying causes and implications for learning dynamics remains preliminary and has not yet been quantitatively established.
Acknowledgments
We thank Luis Iván Hernández Ruíz for helpful discussions. This work was partially supported by JSPS KAKENHI JP22H05106 and JP23H03355 (to HS), JST CREST JPMJCR21N3 (to HS), JSPS KAKENHI JP25K24366 (to HY), and JST BOOST JPMJBS2407 (to YT and MO).
Appendix A Details of Fig. 3
In Fig. 3, the vertical axis, i.e., the KL divergence values, is plotted using a symmetric logarithmic scale (symlog). The scale is linear up to 0.1 bits/byte and logarithmically scaled beyond that point. For the color bar, values are clipped at 10 bits/byte.
Detailed statistics corresponding to Fig. 3 are presented in Table 1(a). In addition, for Pythia-410M, we report the KL divergence between checkpoints separated by 1k steps during pretraining, as well as the KL divergence across different seeds and model sizes at the final checkpoint, in Table 1(b). Furthermore, for the three concrete models, the KL divergence between selected layers is shown in Table 1(c).
Median Mean SD
Later 0.011 0.014 0.0050
Early 0.067 0.078 0.047
Seed 0.12 0.12 0.0071
Size 0.48 0.61 0.47
8-bit 0.44 0.80 1.9
4-bit 0.49 0.91 2.0
FT 0.40 1.5 3.6
Intra-type 0.95 2.8 6.4
Random 2.2 5.3 8.8
Layer 3.0 8.1 13
(a) Summary statistics (median, mean, SD) of KL divergence across various settings. This table corresponds to Fig. 3.
@@ Acknowledgments
(b) KL divergence for Pythia 410M, measured between checkpoints saved during pretraining, across different random seeds, and across different model sizes. For the inter-seed and inter-size comparisons, the final checkpoints are used.
[…]
@@
(c) KL divergence between adjacent layers within specific models, measured using the logit lens.